# Report for NSF Workshop on AI for Electronic Design Automation

### v1

@@ V-D Future Directions and Recommendations to NSF
  • A neurosymbolic symbiotic loop between solvers/provers/verifiers and AI: NSF should promote research in neurosymbolic AI techniques that enable development of ML techniques to improve solvers/provers/verifiers, and in the reverse direction use solvers/provers/verifiers and domain knowledge to improve the trustworthiness of ML.
  • A neurosymbolic symbiotic loop between synthesizers and AI: AI tools, such as LLMs, are surprisingly good at generating code, proofs, and other formal objects. Having said that, with corrective feedback from provers and verifiers during training and fine-tuning (e.g., RLSF) the quality of their output can be dramatically improved. Hence, NSF should focus on combining synthesizers in corrective feedback loops with AI systems.
@@ VI Acknowledgments
This workshop was sponsored by the National Science Foundation, Google, AMD, and The Institute for Learning-enabled Optimization at Scale (TILOS). We thank the speakers and panelists for their insightful contributions, including Deming Chen, Jeff Dean, Vijay Ganesh, Aditya Grover, Farinaz Koushanfar, Koen Lampaert, Yingyan (Celine) Lin, Yong Liu, Igor Markov, Subhasish Mitra, Bryan Perozzi, Pranay Prakash, Ruchir Puri, Mark Ren, Shobha Vasudevan, Yusu Wang, and Lingming Zhang. We also thank the workshop organizers for their leadership and coordination, including Jason Cong, Sergio Guadarrama, Stefanie Jegelka, David Z. Pan, and Yizhou Sun. We gratefully acknowledge the support of the NSF program directors, including Raj Acharya, Sankar Basu, and Sharon Hu. Finally, we thank the student volunteers who assisted with the organization, including Zijian Ding, Weikai Li, Zongyue Qin, and Derek Xu.

### v3

@@ V-D Future Directions and Recommendations to NSF
  • A neurosymbolic symbiotic loop between solvers/provers/verifiers and AI: NSF should promote research in neurosymbolic AI techniques that enable the development of ML techniques to improve solvers/provers/verifiers, and in the reverse direction, use solvers/provers/verifiers and domain knowledge to improve the trustworthiness of ML.
  • A neurosymbolic symbiotic loop between synthesizers and AI: AI tools, such as LLMs, are surprisingly good at generating code, proofs, and other formal objects. Having said that, with corrective feedback from provers and verifiers during training and fine-tuning (e.g., RLSF), the quality of their output can be dramatically improved. Hence, NSF should focus on combining synthesizers in corrective feedback loops with AI systems.
@@ VI Acknowledgments
This workshop was sponsored by the NSF #2436036, Google, AMD, and The Institute for Learning-enabled Optimization at Scale (TILOS). We thank the speakers and panelists for their insightful contributions, including Deming Chen, Jeff Dean, Vijay Ganesh, Aditya Grover, Farinaz Koushanfar, Koen Lampaert, Yingyan (Celine) Lin, Yong Liu, Igor Markov, Subhasish Mitra, Bryan Perozzi, Pranay Prakash, Ruchir Puri, Mark Ren, Shobha Vasudevan, Yusu Wang, and Lingming Zhang. We also thank the workshop organizers for their leadership and coordination, including Jason Cong, Sergio Guadarrama, Stefanie Jegelka, David Z. Pan, and Yizhou Sun. We gratefully acknowledge the support of the NSF program directors, including Raj Acharya, Sankar Basu, and Sharon Hu. Finally, we thank the student volunteers who assisted with the organization, including Zijian Ding, Weikai Li, Zongyue Qin, and Derek Xu.